OpenStudy (anonymous):

Jason has two bags with 6 tiles each. The tiles in each bag are shown below: Without looking, Jason draws a tile from the first bag and then a tile from the second bag. What is the probability of Jason drawing the tile numbered 2 from the first bag and the tile numbered 3 from the second bag? 1 over 36 1 over 12 2 over 12 2 over 6

OpenStudy (muscrat123):

r u trying

OpenStudy (anonymous):

Yes

OpenStudy (muscrat123):

ok good

OpenStudy (anonymous):

@hotguy need ur help

OpenStudy (anonymous):

@muscrat123 i took ur adivce

OpenStudy (muscrat123):

and what is that?

OpenStudy (anonymous):

That openstudyy is not for cheating its for help c:

OpenStudy (muscrat123):

that is exactly right!!! im glad u realized and took my advice

OpenStudy (anonymous):

c: yep and @hotguy u there? :/

OpenStudy (anonymous):

@misssunshinexxoxo HELPPPPP XC

OpenStudy (anonymous):

@muscrat123 would u be able to help :/

OpenStudy (muscrat123):

maybe. need to read it first @wil476003

OpenStudy (anonymous):

okiee

OpenStudy (anonymous):

@manutd30

OpenStudy (anonymous):

ANYONE HELP PLEASEEEEEEEE im on a timed test and need help

OpenStudy (anonymous):

This isn't a graded test right?

OpenStudy (anonymous):

nononon

OpenStudy (anonymous):

Ok, good.

OpenStudy (anonymous):

its for practice

OpenStudy (anonymous):

do u know the answer? by any chance or how

OpenStudy (anonymous):

Well what's the probability of getting a 2 out of the 6 tiles? And what's the probability of getting a 3 out of the 6 tiles?

OpenStudy (anonymous):

Nope, it's not

OpenStudy (anonymous):

hmmmm i know its over 12 doe right

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

no.... :c

OpenStudy (anonymous):

the only possible answer is 1 over 36

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

Multiply the probabilities \[\large \sf \frac{1}{6} \times \frac{1}{6}\]

OpenStudy (anonymous):

2 over 6

OpenStudy (anonymous):

yes? so my answer is 2 over 6

OpenStudy (anonymous):

thats the part i usually get confused on

OpenStudy (anonymous):

\[\large \sf \frac{1}{6} \times \frac{1}{6}~=~\frac{1 \times 1}{6 \times 6}\] @liana1026